# FLASH: Latent-Aware Semi-Autoregressive Speculative Decoding for Multimodal Tasks

### v2

@@ 4.2.1 Video captioning
FLASH $\underline{2.46}$ $\underline{1.83\times}$ FLASH 2.36 $\underline{1.71\times}$ 11.3T
4.2.2 Visual instruction tuning
In visual instruction tuning, each input consists of an image paired with a textual instruction, covering a broad range of tasks such as image captioning, visual question answering, object recognition, and visual reasoning. Using LLaVA-1.5 and QwenVL-2.5 as the target models, we conduct the instruction tuning experiments following the same experiment setups as in the video captioning task.
As presented in Table 2, by comparing the “Text-only” with “Multimodal” speculative decoding, we find that relying solely on text input leads to a significant drop in the draft generation quality when fine-grained image understanding is required, particularly on LLaVA. Additionally, our proposed visual token compression (“VisComp”) and semi-autoregressive (“SemiAR”) brings overall $2.3\%$ and $7.9\%$ speed-up ratio improvements, comparing to “Multimodal”. Specifically, we further find that when temperature $\tau=0$, using “SemiAR” yields a 0.22×0.$2$ speed-up, which is notably higher than the 0.08×0.08$\times$0.08 × improvement observed at $\tau=1$. This improvement due to the model’s tendency to produce more deterministic outputs at lower temperatures, which aligns well with the parallel token generation in the semi-autoregressive head.
Notably, “SemiAR” yields greater improvements in visual instruction tuning than in video captioning. Besides, we observe that the improvement brought by visual token compression is less pronounced in the visual instruction tuning task compared to the video captioning task. This is likely because visual instruction tuning typically involves a single image, whereas video captioning tasks process multiple frames. By combining “VisComp” and “SemiAR”, FLASH achieves speed-up ratios of 2.55×2.$5$ and 1.83×1.83$\times$1.83 × on LLaVA and QwenVL, respectively. Moreover, we observe that FLASH efficiently processes visual information while incurring a computational load comparable to that of the “Text-only” model. It enables FLASH to achieve a significantly higher speed-up ratio without sacrificing draft quality. These results further demonstrate the effectiveness of our method across both video captioning and visual instruction tuning tasks.
4.3 Ablations
@@
5 Conclusion and Limitations
In this work, we introduce FLASH, a novel speculative decoding framework designed specifically for Large Multimodal Models (LMMs). FLASH introduces two key components: a visual token compression module that reduces input redundancy, and a semi-autoregressive decoding strategy that enables parallel token generation. Together, these innovations accelerate multimodal inference while preserving the quality of the output. Our experiments show that FLASH achieves substantial speed-up over prior speculative decoding methods across video captioning and visual instruction tuning tasks. These results demonstrate the potential of FLASH to enable more efficient deployment of LMMs.
For video tasks, we employ a sampling method where one frame is selected per second, following previous work Zhang et al. (2025). However, it may not capture and preserve the dynamic information in videos. It is important to note that the primary focus of this work is on training a draft model to predict the generation of the target LMM, rather than optimizing temporal dynamics. Future work will extend this framework to support more multimodal tasks and explore more strategies, such as draft length prediction to further improve the draft quality.
Appendix A Appendix
A.1 Scalability
To validate the robustness of our approach, we further conducted experiments using LLaVA-13B, QwenVL-3B,QwenVL-32B, in addition to the main results reported on LLaVA-7B and QwenVL-7B. The results in Table 3 and Table 4 show that FLASH consistently achieves a higher speed-up ratio compared to multimodal speculative decoding, a variant of Eagle adapted to handle multimodal input. This highlights the efficiency of FLASH in both video captioning and instruction tuning tasks.
@@ A.1 Scalability
Multimodal 1.63×1.63$\times$1.63 × 1.24×1.24$\times$1.24 × 2.48×2.48$\times$2.48 × 2.49×2.49$\times$2.49 × 2.01×2.01$\times$2.01 ×
FLASH 1.83×1.83$\times$1.83 × 1.47×1.47$\times$1.47 × 2.65×2.65$\times$2.65 × 2.68×2.68$\times$2.68 × 2.20×2.20$\times$2.20 ×
Table 4: Quantitative experiments of instruction tuning, reporting the speed-up ratio as the primary metric.
@@
Multimodal 2.21×2.21$\times$2.21 × 1.21×1.21$\times$1.21 × 1.88×1.$8$ 1.63×1.63$\times$1.63 × 1.40×1.40$\times$1.40 ×
FLASH 2.55×2.$5$ 1.49×1.49$\times$1.49 × 2.01×2.01$\times$2.01 × 1.83×1.83$\times$1.83 × 1.49×1.49$\times$1.49 ×

### v3

@@ Appendix B Algorithm
5: // Visual token compression
[…]
7: // Semi-autoregressive generation
[…]
12:Return $T^{\prime}_{i},\dots,T^{\prime}_{K}\sim\text{Categorical}(P^{\prime}_{1}),\dots,\text{Categorical}(P^{\prime}_{K})$
Appendix C Ablations on Compression Ratios
We further conduct ablations on the scale of $\mathcal{C}$, which influences the number of visual tokens after compression. The results are presented in Table 7. Our chosen ratio of $1/9$ strikes an optimal balance, delivering the peak speed-up. The performance remains robust across a range of ratios (from $1/4$ to $1/16$), indicating that SpecFLASH is not overly sensitive to the exact compression rate. However, the extreme case of removing visual tokens entirely ($C/N=0$) leads to a clear performance drop, confirming that our compression module retains essential visual information necessary for maintaining draft quality.
Table 7: Speed-up Ratios with different compression ratios on visual instruct tuning (VIT) and video captioning (VC).
Compression ratios C/N VIT VC
[…]
Appendix D Ablation on Draft Length K
@@
Appendix E Ablation on Semi-autoregressive Group Size $k^{\prime}$
We further conduct ablations on the number of draft tokens produced per forward pass of the draft model $k^{\prime}$. For example, when semi-autoregressive group size is $k^{\prime}=2$ and the draft length $K=4$, the draft model first generates 2 draft tokens, then uses these tokens as input to generate the next 2 draft tokens; the resulting 4 draft tokens are subsequently evaluated by the target model. The results in Table 8 show that setting $k^{\prime}=4$ yields better speed-up ratios. In this case, the semi-autoregressive head predicts 4 tokens in a single forward pass, and these candidate tokens are then evaluated by the target model.
Table 8: Speed-up ratios with different $k^{\prime}$ settings.
Method $k^{\prime}$ $K$ Speed-up ratio
SpecFLASH (main results) 4 4 $2.55\times$
SpecFLASH 2 4 $2.47\times$
Autoregressive 1 4 $2.23\times$
@@ Appendix E Ablation on Semi-autoregressive Group Size $k^{\prime}$
Eagle-$M$ $1.63\times$ $1.24\times$ $2.48\times$ $2.49\times$ $2.01\times$
SpecFLASH $1.83\times$ $1.47\times$ $2.65\times$ $68\times 2.2$$20\times$
Table 10: Quantitative experiments of visual instruction tuning, reporting the speed-up ratio as the primary metric.
@@
Eagle-$M$ $2.21\times$ $1.21\times$ $1.88\times$ $1.63\times$ $1.40\times$
SpecFLASH $2.55\times$ $1.49\times$ $2.01\times$ $1.83\times$ $1.49\times$
Appendix F Scalability
To validate the robustness of our approach, we further conducted experiments using LLaVA-13B, QwenVL-3B,QwenVL-32B, in addition to the main results reported on LLaVA-7B and QwenVL-7B. The results in Table 9 and Table 10 show that SpecFLASH consistently achieves a higher speed-up ratio compared to multimodal speculative decoding, a variant of Eagle adapted to handle multimodal input (Eagle-MM). This highlights the efficiency of SpecFLASH in both video captioning and instruction tuning tasks.
Appendix G Multi-Task Evaluation of Visual Instruction Tuning
Visual instruction tuning encompasses multiple tasks, such as fine-grained image captioning, ScienceQA, and commonsense reasoning. We evaluate the performance on each benchmark individually. The results demonstrate that SpecFLASH consistently outperforms the competing methods across all tasks, indicating its robustness.
Table 11: Performance comparison on multiple tasks.
Method COCO Captioning MMT-Bench ScienceQA
EAGLE-$M$ $28\times 2.2$$26\times$ $2.15\times$
DREAM $49\times 2.2$$23\times$ $2.11\times$
SpecFLASH $2.64\times$ $2.38\times$ $2.32\times$
Appendix H Limitations and Future works
Experiments demonstrate that SpecFLASH consistently accelerates both video captioning and visual instruction tuning. Importantly, SpecFLASH has the potential to serve as a general-purpose acceleration method on a wide range of multimodal tasks, as visual instruction tuning itself encompasses multiple such tasks. We intend to extend our approach to additional multimodal tasks in future work. Furthermore, the semi-autoregressive decoding strategy can experience reduced acceptance rates as the draft length increases. Although these trade-offs did not significantly impact the results reported here, they indicate that further investigation into adaptive compression techniques and more flexible decoding strategies will be important for achieving broader and more reliable deployment.